Kim Inc. must install a new air conditioning unit in its main plant. Kim must install one or the other of the units; otherwise, the highly profitable plant would have to shut down. Two units are available, HCC and LCC (for high and low capital costs, respectively). HCC has a high capital cost but relatively low operating costs, while LCC has a low capital cost but higher operating costs because it uses more electricity. The costs of the units are shown here. Kim's WACC is 7.5%. 0 1 2 3 4 5 HCC -$590,000 -$45,000 -$45,000 -$45,000 -$45,000 -$45,000 LCC -$90,000 -$170,000 -$170,000 -$170,000 -$170,000 -$170,000 Which unit would you recommend?
Answer #1

NPV of HCC=-590000-45000/7.5%*(1-1/1.075^5)=-772064.8206

NPV of LCC=-90000-170000/7.5%*(1-1/1.075^5)=-777800.4333

Since we are examining costs, the unit chosen would be the one that had the lower NPV of costs. Since HCC's NPV of costs is lower than LCC's, HCC would be chosen.